# FLARE: Faithful Logic-Aided Reasoning and Exploration

### v2

@@ 2.2 Detecting Reasoning Inconsistencies
With this framework in mind, we define two reasoning failure modes. In the first failure mode, given that some fact or relation was used in the simulated search but did not exist in the generated code, i.e. $\exists j\text{ such that }\mathcal{F}_{\text{search}}^{j}\notin\mathcal{F}_{\text{code}}$, we claim that the LLM has hallucinated. We postulate that the model either produced incomplete knowledge during formalisation to code or created a piece of non-existing information during the search. We do not consider facts that emerged during a direct inference step within the simulated search during our calculation. For example, if we are dealing with a mathematical query $4\cdot(5+6)=?$, the search would involve separately evaluating the expression $5+6=11$. In this case, $11$ will not be treated as a hallucinated fact within the search but rather as an emergent fact obtained from direct inference. The second failure mode is the reciprocal case, where a fact or relation present in the code is not used during the search. We refer to this phenomenon as sub-optimal reasoning as it shows that the LLM could not explore the problem space completely or injected unsuitable knowledge during formalisation into code.
[…]
@@ 3.1 Datasets
To evaluate FLARE, we use a benchmark of $9$ tasks that cover Math Word Problems (MWP), multi-hop QA and relation inference. For testing numeric and mathematical reasoning, we follow CoT (Wei et al., 2022) by including GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), MultiArith (Roy & Roth, 2015), ASDiv (Miao et al., 2020) and AQuA (Ling et al., 2017). Among these, GSM8K, SVAMP, MultiArith and ASDiv cover elementary and middle school arithmetic word problems with a set of integers or decimals as the answer. AQuA is a multiple-choice numerical, symbolic reasoning dataset where each answer is a mathematical expression containing notations, values and expressions not defined in the query. We also test FLARE using three multi-hop QA datasets. We use StrategyQA (Geva et al., 2021), which is a boolean QA task that requires sub-goal decomposition and a multi-hop reasoning strategy to answer. The example “Do all parts of the aloe vera plant taste good?” used in Figure 1, is taken from StratedyQA. The multi-hop QA testing also includes Date and Sports Understanding, subsets of BIG-Bench (bench authors, 2023). The tasks involve inferring an exact date given some calculations in the relative time period and understanding if an artificially created sports statement is feasible. Finally, we assess FLARE on Relational Inference using CLUTRR (Sinha et al., 2019), which involves inferring the familial relation between two entities mentioned in a natural language description of the partial family graph. We adopt the same in-context samples used in (Xu et al., 2024; Wei et al., 2022). The complete descriptions, statistics and examples for each dataset can be found in Table 7 in Appendix A.
Figure 2: The trend of mean model accuracy w.r.t mean faithfulness (ROUGE-Lsum) for all the models. Faithfulness is positively correlated with model performance.
@@ 3.2 Benchmarks
We compare FLARE with CoT (Wei et al., 2022) as a prompting method that reasons using natural language chains and with F-CoT (Xu et al., 2024) that formalises the query into a code and offload the reasoning to an external symbolic solver. We use Llama3.1 (8B) (Dubey et al., 2024), CmDR (30B) (Cohere, 2024), CmDR+ (100B) (Cohere, 2024) and GPT3.5 (Brown et al., 2020b) ($\geq 100B$ (Ye et al., 2023)). As the coding model OpenAI Codex (code-DaVinci-002) (Chen et al., 2021) used in F-CoT has been deprecated, we replace it with the new GPT3.5 as suggested by OpenAI and recalculate the results accordingly.
@@ 4.1 Main Results
To evaluate FLARE, we use a set of models of varying sizes on diverse benchmarks, as defined in Section 3. We compare the performance of each model while using FLARE, CoT and F-CoT prompting. The results for F-CoT and CoT on all the models are computed using the codebase of the original study (Xu et al., 2024).
@@
Our results, presented in Table 1, show that using FLARE allows the LLMs to achieve state-of-the-art results on $7$ out of $9$ datasets, with an average $28\%$ increase over CoT. We can see a clear trend that FLARE increases the performance compared to CoT and F-CoT for all the models of varying scales. We also see that LLMs that are not explicitly tuned for coding suffer massive degeneracies when using F-CoT. We postulate that they are unable to consistently produce executable programs that satisfy a predefined scheme in F-CoT, thus resulting in an error during execution. This further highlights the value of simulating program execution using an LLM instead of using external solvers. The results show that using FLARE yields more benefit on datasets that require longer chains of multi-hop and symbolic reasoning, like AQuA and StrategyQA.
@@ 4.4 What is important during the search?
[…]
We expand the analysis of the simulated search traces to detect the reasons which can lead to optimal reasoning within an LLM. For this purpose, we calculate several statistics, like the average number of explored paths, average and total hops and failures per path, for each model during the simulated traversal. The failure in a path is an invalidation of a solution for a sub-goal explored during the search, which is used for backtracking, as explained in Section 2. Calculating these statistics is simple as the search component of FLARE, seen in Figure 1, is a structured simulation of a Prolog trace, where each line contains a hop of reasoning inference. We split these statistics for the reasoning paths that lead to correct or incorrect outcomes. Our results in Table 3 show that LLM performance and reasoning optimally are not directly connected to the amount of explored paths or multi-hop inferences per path. We also see that traces that lead to incorrect answers have a higher number of failures per path and in total. We explain this phenomenon with the hypothesis that LLMs with traces that were optimal for reasoning and led to correct answers could skip exploring degenerate solutions due to strong commonsense reasoning capabilities. Further analyses focus on identifying inconsistencies and failure modes (Section 2.2). By comparing relations in code with those in search traces, we measure emergent hallucinations and unused relations, highlighting areas of sub-optimal reasoning. Additionally, we assess the uniqueness of emergent facts per inference hop, which indicates the extent of problem-space exploration (Table 4). The results in Table 4 show consistently over each model that, on average, traces that lead to correct answers had a higher percentage of unique emergent facts and overlap in the relations used between the code and search, while the portion of underutilized relations was lower. This means that optimal reasoning with an LLM requires a great degree of problem-space exploration with fewer relation hallucinations during the search and more relation utilization from the defined code. This aligns with our prior discoveries, which show a strong correlation between simulated search faithfulness towards the formalised code and model performance. Our framework FLARE has these reasoning patterns ingrained within its inference pipeline.
@@ 4.5 The effect of scale
5 Related Work
Reasoning in Natural Language
Reasoning with Search
Several lines of work propose using techniques to expand the reasoning paths over the problem space. Self-consistency decoding (Wang et al., 2023b) is an approach used to sample many natural language reasoning paths and take a majority vote for an answer. Another popular approach is Tree-of-Thoughts (ToT, Yao et al. (2023a)), which proposes to explore the problem space with reasoning similar to a tree traversal, where each state is created and evaluated using an LLM. Similar techniques try to adapt symbolic search approaches akin to DFS, BFS (Besta et al., 2024), A (Lehnert et al., 2024) or other combinations (Gandhi et al., 2024) with direct tuning (Lehnert et al., 2024), imitation training (Yang et al., 2022) or few-shot prompting (Zhang et al., 2024). It must be noted that all of these techniques have only been tested in constrained mathematical puzzle-solving and algorithmic domains like the 24 Game (Yang et al., 2022), Countdown (Wikipedia, 2024), Sorting (Besta et al., 2024), maze solving (Yang et al., 2022), Sokoban (Lehnert et al., 2024) and others. Although the search component of FLARE has some similarities to these techniques, we argue that our method allows for generalistic reasoning with interpretable multi-hop search through iterative logic-based problem space exploration.
Reasoning with Formalisation
Another line of research has tried formalising natural language queries into code (Gao et al., 2023; Li et al., 2024) or pseudo-code (Chae et al., 2024; Gandhi et al., 2024). This allows the translation of the query into a strict structure and offloads the reasoning and search components to deterministic solvers like Python Chen et al. (2023), PDDL Xu et al. (2024); Liu et al. (2023), DataLog Xu et al. (2024) and others. While models are reasonably capable of synthesising programs (Austin et al., 2021; Nijkamp et al., 2023) and benefit from the use of code in numerical and algorithmic reasoning settings (Chen et al., 2023; Gao et al., 2023), the usage of code for general QA has not been rigorously explored. The reasons are that formalisation from natural language into a strict and executable code is challenging (Wu et al., 2022), following the exact syntactic constraints of the programming language not abundantly used during pre-training is onerous (Liu et al., 2024) and can require models explicitly tuned for coding (Chen et al., 2021). Using an external solver for reasoning also limits the capability for soft reasoning in commonsense knowledge and implications. Although we formalise the natural language query into a logic programming Prolog program during the code generation part of FLARE, we do not explicitly require the code to be executable and do not use external solvers during inference. This allows for the further use of the LLM for soft-reasoning to simulate code execution in a logic-based problem space traversal similar to Prolog while circumventing the need for code tuning a generalist model.
Reasoning Faithfulness
@@ 6 Conclusion
Acknowledgments
[Uncaptioned image][Uncaptioned $\end{array}$ Erik is partially funded by a DFF Sapere Aude research leader grant under grant agreement No 0171-00034B, as well as by a NEC PhD fellowship. Pasquale was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. Isabelle’s research is partially funded by the European Union (ERC, ExplainYourself, 101077481). This work was further supported by by the Pioneer Centre for AI, DNRF grant number P1, the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.

### v3

@@ 3.2 Detecting Reasoning Inconsistencies
With this framework in mind, we define two reasoning failure modes.
Dataset ChatGPT (gpt-3.5-turbo) Standard CoT Logic-LM FLARE PrOntoQA 47.40 67.80 61.00 73.40 LogicalDeduction 40.00 42.33 65.67 58.60 AR-LSAT 20.34 17.31 26.41 27.39
Table 2: Comparison of Direct Prompting, CoT, Logic-LM and FLARE.
In the first failure mode, given that some fact or relation was used in the simulated search but did not exist in the generated code, i.e. $\exists j\text{ such that }\mathcal{F}_{\text{search}}^{j}\notin\mathcal{F}_{\text{code}}$, we claim that the LLM has hallucinated. We postulate that the model either produced incomplete knowledge during formalisation to code or created a piece of non-existing information during the search. We do not consider facts that emerged during a direct inference step within the simulated search during our calculation. For example, if we are dealing with a mathematical query $4\cdot(5+6)=?$, the search would involve separately evaluating the expression $5+6=11$. In this case, $11$ will not be treated as a hallucinated fact within the search but rather as an emergent fact obtained from direct inference. The second failure mode is the reciprocal case, where a fact or relation present in the code is not used during the search. We refer to this phenomenon as sub-optimal reasoning as it shows that the LLM could not explore the problem space completely or injected unsuitable knowledge during formalisation into code.
@@ 4.1 Datasets
To evaluate FLARE, we use a benchmark of $9$ tasks covering Math Word Problems (MWP), multi-hop QA and relation inference, and $3$ common logical reasoning datasets. For testing numeric and mathematical reasoning, we follow CoT (Wei et al., 2022) by including GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), MultiArith (Roy & Roth, 2015), ASDiv (Miao et al., 2020) and AQuA (Ling et al., 2017). Among these, GSM8K, SVAMP, MultiArith and ASDiv cover elementary and middle school arithmetic word problems with a set of integers or decimals as the answer. AQuA is a multiple-choice numerical, symbolic reasoning dataset where each answer is a mathematical expression containing notations, values and expressions not defined in the query. We also test FLARE using three multi-hop QA datasets. We use StrategyQA (Geva et al., 2021), which is a boolean QA task that requires sub-goal decomposition and a multi-hop reasoning strategy to answer. The example “Do all parts of the aloe vera plant taste good?” used in Figure 1, is taken from StrategyQA. The multi-hop QA testing also includes Date and Sports Understanding, subsets of BIG-Bench (bench authors, 2023). The tasks involve inferring an exact date given some calculations in the relative time period and understanding if an artificially created sports statement is feasible. Furthermore, we assess FLARE on Relational Inference using CLUTRR (Sinha et al., 2019), which involves inferring the familial relation between two entities mentioned in a natural language description of the partial family graph. We evaluate FLARE on challenging logic inference datasets: ProntoQA (Saparov & He, 2023), AR-LSAT (Zhong et al., 2021), and LogicalDeductions from BigBench (et al., 2023), focusing on harder subsets proposed by Pan et al. (2023). These datasets, covering deductive, analytical, and logical reasoning, allow us to assess FLARE’s performance. Details, including descriptions and examples, are in Table 9 of Appendix A.
@@ 4.2 Benchmarks
We compare FLARE with CoT (Wei et al., 2022) as a prompting method that reasons using natural language chains and with F-CoT (Lyu et al., 2023) and Logic-LM (Pan et al., 2023) that formalise the query into a code and offload the reasoning to an external symbolic solver. We use Llama3.1 (8B) (Dubey et al., 2024), CmDR (30B) (Cohere, 2024), CmDR+ (100B) (Cohere, 2024) and GPT3.5 (Brown et al., 2020b) ($\geq 100B$ (Ye et al., 2023)). As the coding model OpenAI Codex (code-DaVinci-002) (Chen et al., 2021) used in F-CoT has been deprecated, we replace it with the new GPT3.5 as suggested by OpenAI and recalculate the results accordingly.
@@ 5.1 Few-shot prompting
To evaluate FLARE, we use a set of models of varying sizes on diverse benchmarks, as defined in Section 4. We compare the performance of each model while using FLARE, CoT and F-CoT prompting. The results for F-CoT and CoT on all the models are computed using the codebase of the original study (Lyu et al., 2023). We additionally compare Logic-LM and FLARE using the logic reasoning benchmarks proposed in Pan et al. (2023).
@@
Our results, presented in Table 1, show that using FLARE allows the LLMs to achieve state-of-the-art results on $7$ out of $9$ datasets, with an average $28\%$ increase over CoT. We can see a clear trend that FLARE increases the performance compared to CoT and F-CoT for all the models of varying scales. We also see that LLMs that are not explicitly tuned for coding suffer massive degeneracies when using F-CoT. We postulate that they are unable to consistently produce executable programs that satisfy a predefined scheme in F-CoT, thus resulting in an error during execution. This further highlights the value of simulating program execution using an LLM instead of using external solvers. The results show that using FLARE yields more benefit on datasets that require longer chains of multi-hop and symbolic reasoning, like AQuA and StrategyQA. Our findings in Table 2 show that FLAREachieves state-of-the-art results on $2$ out of $3$ logic inference benchmarks with $10\%$ increase over CoT and ${}7\%$ increase over Logic-LM.
@@ 5.4 What is important during the search?
We expand the analysis of the simulated search traces to detect the reasons which can lead to optimal reasoning within an LLM. For this purpose, we calculate several statistics, like the average number of explored paths, average and total hops and failures per path, for each model during the simulated traversal. The failure in a path is an invalidation of a solution for a sub-goal explored during the search, which is used for backtracking, as explained in Section 3. Calculating these statistics is simple as the search component of FLARE, seen in Figure 1, is a structured simulation of a Prolog trace, where each line contains a hop of reasoning inference. We split these statistics for the reasoning paths that lead to correct or incorrect outcomes. Our results in Table 4 show that LLM performance and reasoning optimally are not directly connected to the amount of explored paths or multi-hop inferences per path. We also see that traces that lead to incorrect answers have a higher number of failures per path and in total. We explain this phenomenon with the hypothesis that LLMs with traces that were optimal for reasoning and led to correct answers could skip exploring degenerate solutions due to strong commonsense reasoning capabilities. Further analyses focus on identifying inconsistencies and failure modes (Section 3.2). By comparing relations in code with those in search traces, we measure emergent hallucinations and unused relations, highlighting areas of sub-optimal reasoning. Additionally, we assess the uniqueness of emergent facts per inference hop, which indicates the extent of problem-space exploration (Table 5). The results in Table 5 show consistently over each model that, on average, traces that lead to correct answers had a higher percentage of unique emergent facts (UEF) and overlap in the relations (OR) used between the code and search, while the portion of underutilized relations (UR) was lower. This means that optimal reasoning with an LLM requires a great degree of problem-space exploration with fewer relation hallucinations during the search and more relation utilization from the defined code. This aligns with our prior discoveries, which show a strong correlation between simulated search faithfulness towards the formalised code and model performance. Our framework FLARE has these reasoning patterns ingrained within its inference pipeline.
@@ A.2 Dataset Statistics
A.3 FLARE Pseudo-code
Below, we present the pseudo-code for the execution of the plan, code, and search procedures in FLARE. The pseudo-code describes the modular pipeline in FLARE for tackling natural language queries with faithful simulated search.
  • Plan Generation: This stage creates a structured natural language outline of the reasoning process, breaking down the query into logical steps and analysis. The plan serves as the foundation for formalization into a logic-based representation.
  • Code Generation: Based on the generated plan, a logic programming code (e.g., in Prolog) is synthesized. This code formalizes the query into a set of facts, relations, and goals, which collectively define the problem space for reasoning.
  • Search Simulation: The generated code is utilized to simulate a search trace over the problem space. This includes iterative reasoning, backtracking when goals are unmet, and extracting emergent facts or relations during the process.
Each of these stages is implemented as a modular component. The generation from each of the stages feeds into the next, allowing seamless integration and incremental improvement in reasoning accuracy. A detailed pseudo-code is provided below in Algorithm 1.
Algorithm 1 FLARE Methodology: Faithful Logic-Aided Reasoning and Exploration
1:Query $\mathcal{Q}$, Language Model $\mathcal{M}$
2:Answer $\mathcal{A}$
3:Initialization: Load few-shot examples for plans ($\mathcal{E}_{P}$), code ($\mathcal{E}_{C}$), and search traces ($\mathcal{E}_{S}$)
4:Input: Natural language query $\mathcal{Q}$
5:procedure Generate Plan
6:     Prompt $\mathcal{M}$ with instructions $\mathcal{I}_{P}$ and examples $\mathcal{E}_{P}$ to generate a plan $\mathcal{P}$
7:     $\mathcal{P}\leftarrow\arg\max\;p_{\mathcal{M}}(T_{P}|T_{P:<i},\mathcal{E}_{P},\mathcal{Q},\mathcal{I}_{P})$
8:end procedure
9:procedure Generate Code
10:     Append examples $\mathcal{E}_{C}$ to $\mathcal{E}_{P}$
11:     Prompt $\mathcal{M}$ with instructions $\mathcal{I}_{C}$ to generate logic programming code $\mathcal{C}$
12:     $\mathcal{C}\leftarrow\arg\max\;p_{\mathcal{M}}(T_{C}|T_{C:<i},\mathcal{E}_{C},\mathcal{Q},\mathcal{P},\mathcal{I}_{C})$
13:     $(F_{\text{code}},R_{\text{code}},G_{\text{code}})\leftarrow\textsc{Extract}(\mathcal{C})$
14:end procedure
15:procedure Simulate Search
16:     Append search trace examples $\mathcal{E}_{S}$ to $\mathcal{E}_{C}$
17:     Prompt $\mathcal{M}$ with instructions $\mathcal{I}_{S}$ to simulate a search trace $\mathcal{S}$
18:     $\mathcal{S}\leftarrow\arg\max\;p_{\mathcal{M}}(T_{S}|T_{S:<i},\mathcal{E}_{S},\mathcal{Q},\mathcal{P},\mathcal{C},\mathcal{I}_{S})$
19:     $(F_{\text{search}},R_{\text{search}},\mathcal{A}_{\text{search}})\leftarrow\textsc{Extract}(\mathcal{S})$
20:     while Goal $\mathcal{G}_{\text{code}}$ is not satisfied do
21:         Explore next sub-goal in $\mathcal{S}$
22:         if Sub-goal fails then
23:              Backtrack to the previous state (Learned through in-context sampels)
24:         end if
25:     end while
26:end procedure
27:procedure Final Answer Generation
28:     Append correct answers from $\mathcal{A}_{\text{search}}$ to examples
29:     Prompt $\mathcal{M}$ with instructions $\mathcal{I}_{A}$ to finalize answer $\mathcal{A}$
30:     $\mathcal{A}\leftarrow\arg\max\;p_{\mathcal{M}}(T_{A}|T_{A:<i},\mathcal{E}_{A},\mathcal{Q},\mathcal{P},\mathcal{C},\mathcal{S},\mathcal{I}_{A})$
31:end procedure
32:return $\mathcal{A}$
Section Content
Question ”Can Reiki be stored in a bottle?”
Explanation Reiki is a form of alternative therapy that involves the channeling of energy to promote healing and relaxation. It is often administered through the practitioner’s hands, with the energy flowing directly to the recipient. The question raises the possibility of storing Reiki energy in a bottle, which may be a common query due to the abstract nature of energy healing.
Analysis To address the question of whether Reiki can be stored in a bottle, it is important to consider the nature of Reiki energy, the principles of energy healing, and the practical implications of containing and preserving such energy.
Plan Understand Reiki Energy: … Explore Energy Storage: … Consider Symbolism and Intention: … Evaluate Practicality: … Reflect on Tradition: …
Code % Define the nature of Reiki energy reiki_energy(spiritual_energy). reiki_energy(channeling_through_touch). % Define the feasibility of storing energy in a bottle store_energy_in_bottle(feasible) :- reiki_energy(spiritual_energy), reiki_energy(channeling_through_touch). % Define the assumption that storing Reiki energy in a bottle is not a common practice not_common_practice(store_reiki_in_bottle). % Query to determine if Reiki can be stored in a bottle query :- store_energy_in_bottle(feasible), not_common_practice(store_reiki_in_bottle).
Search Paths Path 1: 0: Start of execution: Beginning Search 1: Call: query 2: Call: store_energy_in_bottle(feasible) 3: Call: reiki_energy(spiritual_energy) 4: Call: reiki_energy(channeling_through_touch) 5: Call: not_common_practice(store_reiki_in_bottle) 6: Call: feasible=no 7: Fail: feasible=no — ’Result’: ’Search Failed’ 8: Fail: query — ’Result’: ’Search Failed’
Answer ”No, Reiki energy cannot be stored in a bottle based on the logical evaluation of its abstract, non-physical nature and traditional practices of Reiki.”
Table 8: Complete example of FLARE